# SPLITZ: Certifiable Robustness via Split Lipschitz Randomized Smoothing thanks: This work was supported by NSF grants CCF 2100013, CNS 2209951, CCF 1651492, CNS 2317192 and CNS 1822071. This wrok has been submitted to IEEE Transactions on Information Forensics and Security

### v1

@@ V Discussion and conclusion
In this paper, we presented SPLITZ, a novel and practical certified defense mechanism, where we constrained the local Lipschitz bound of the left half of the classifier and smoothed the right half of the classifier with noise. This is because the local Lipschitz constant can capture information specific to each individual input, and the relative stability of the model around that input. To the best of our knowledge, this is the first systematic framework to combine the ideas of local Lipschitz constants with randomized smoothing. Furthermore, we provide a closed-form expression for the certified radius based on the local Lipschitz constant of the left half of the classifier and the randomized smoothing based radius of the right half of the classifier. We show that maintaining a relatively small local Lipschitz constant of the left half of the classifier helps to improve the certified robustness (radius). We showed results on several benchmark datasets and obtained significant improvements over state-of-art methods for MNIST, CIFAR-10 datasets. For instance, on the CIFAR-10 dataset, SPLITZ can achieve $43.2\%$ certified test accuracy compared to state-of-art certified test accuracy $39.8\%$ with $\ell_{2}$ norm perturbation budget of $\epsilon=1$. We reported similar results for the MNIST dataset. We believe that combining the core idea of SPLITZ with other recent techniques, such as denoising diffusion models and adversarial training can be a fruitful next step to further improve certified robustness.
@@
Appendix A Additional Results on Lipschitz bounds
@@ Appendix B Lipschitz Constrained Training
$D^{v}(i,i)=\begin{cases}\mathbbm{1}(\text{ReLU}\theta(z_{l}^{i})>0),&\text{if}~{}~{}I^{v}(i,i)=1\\ 0,&\text{otherwise}\\ \end{cases}$ (23)
@@
$L_{local}(x,f)\leq\parallel W_{L}I^{v}_{L-1}\parallel\parallel I^{v}_{L-1}W_{L-1}I^{v}_{L-2}\parallel\dots\parallel I^{v}_{1}W_{1}\parallel$ (24)
As stated in [19], it straight forward to prove $\parallel I^{v}_{L-1}W_{L-1}I^{v}_{L-2}\parallel\leq\parallel W_{L-1}\parallel$ using the property of eigenvalues. We briefly prove it following from [19].
@@
where $x\in\mathcal{R}^{3}$ and $y\in\mathcal{R}$ and $W^{l}$ denotes the weight matrix for layer $l$. Moreover the threshold $\theta=1$.
@@ Appendix C Additional Experimental Details
In this section, we provide additional results for the two datasets, e.g., MNIST, CIFAR-10 dataset. We first provide the details of two datasets. Next, we illustrate the baselines used in our paper. Note that we report the numbers (certified test accuracy, average certified radius) from respective papers.
C-A Training details
For all value of $\sigma$, we keep the value of training $\sigma$ and testing $\sigma$ to be the same. We apply the noise samples $n_{0}=100$ to predict the most probably class $c_{A}$ and denote $\alpha=0.001$ as the confidence during the certifying process. Furthermore, we use $n_{1}=100,000$ to calculate the lower bound of the probability $p_{A}$ for the MNIST, CIFAR-10 respectively. Moreover, to maintain a relatively small local Lipschitz constant of left half of the SPLITZ classifier, we set the threshold of clipped ReLU (see Sec B) as 1 for all datasets. For estimating the local Lipschitz constant of the left half of the classifier, the power iteration is 5, 2 during the training for MNIST, CIFAR-10 respectively following from [19].
C-B Details of datasets
MNIST [36] dataset contains handwritten digits usually used for image classification problems. The dataset is comprised of a total of 70,000 images, with 60,000 images in the training set and 10,000 in the test set. The dataset has 10 classes and is in grayscale format. We pre-process the MNIST dataset using normalization. We incorporate the Lipschitz constant of the normalization layer when calculating the local Lipschitz constant of the left half of the classifier.
CIFAR-10 [37] dataset consists of 60,000 RGB images distributed across 10 categories: airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck. Each category is represented by 6,000 images. The dataset is divided into a training set with 50,000 images and a test set containing 10,000 images. We employ the conventional data augmentation techniques of random horizontal flipping and random translation by up to 4 pixels, consistent with methods used in other baseline studies [14, 7, 12]. Additionally, we normalize each image on a pixel-by-pixel basis. We incorporate the Lipschitz constant of the normalization layer when calculating the local Lipschitz constant of the left half of the classifier.

### v2

@@ V Discussion and conclusion
In this paper, we presented SPLITZ, a novel and practical certified defense mechanism, where we constrained the local Lipschitz bound of the left half of the classifier and smoothed the right half of the classifier with noise. This is because the local Lipschitz constant can capture information specific to each individual input, and the relative stability of the model around that input. To the best of our knowledge, this is the first systematic framework to combine the ideas of local Lipschitz constants with randomized smoothing. Furthermore, we provide a closed-form expression for the certified radius based on the local Lipschitz constant of the left half of the classifier and the randomized smoothing based radius of the right half of the classifier. We show that maintaining a relatively small local Lipschitz constant of the left half of the classifier helps to improve the certified robustness (radius). We showed results on several benchmark datasets and obtained significant improvements over state-of-the-art methods for the MNIST, CIFAR-10 and ImageNet datasets. For instance, on the CIFAR-10 dataset, SPLITZ can achieve $43.2\%$ certified test accuracy compared to state-of-art certified test accuracy $39.8\%$ with $\ell_{2}$ norm perturbation budget of $\epsilon=1$. We observed similar trends for the MNIST and ImageNet dataset. We believe that combining the core idea of SPLITZ with other recent techniques, such as denoising diffusion models and adversarial training can be a fruitful next step to further improve certified robustness.
@@
Appendix A Additional Lipschitz Constants Results
@@ Appendix B Lipschitz Constrained Training
$D^{v}(i,i)=\begin{cases}\mathbbm{1}(\text{ReLU}\theta(z_{m}^{i})>0),&\text{if}~{}~{}I^{v}(i,i)=1\\ 0,&\text{otherwise}\\ \end{cases}$ (23)
@@
$L_{local}(x,f)\leq\parallel W_{M}I^{v}_{M-1}\parallel\parallel I^{v}_{M-1}W_{M-1}I^{v}_{M-2}\parallel\dots\parallel I^{v}_{1}W_{1}\parallel$ (24)
As stated in [19], it straight forward to prove $\parallel I^{v}_{M-1}W_{M-1}I^{v}_{M-2}\parallel\leq\parallel W_{M-1}\parallel$ using the property of eigenvalues. We briefly prove it following from [19].
@@
where $x\in\mathcal{R}^{3}$ and $y\in\mathcal{R}$ and $W^{m}$ denotes the weight matrix for layer $m$. Moreover the threshold $\theta=1$.
@@ Appendix C Additional Experimental Details
In this section, we provide additional results for the three datasets, e.g., MNIST, CIFAR-10 and ImageNet dataset. We first provide the details of three datasets. Next, we illustrate the baselines used in our paper. Note that we report the numbers (certified test accuracy, average certified radius) from respective papers.
Training details For all value of $\sigma$, we keep the value of training $\sigma$ and testing $\sigma$ to be the same. We apply the noise samples $n_{0}=100$ to predict the most probably class $c_{A}$ and denote $\alpha=0.001$ as the confidence during the certifying process. Furthermore, we use $n_{1}=100000,100000,10000$ to calculate the lower bound of the probability $p_{A}$ for the MNIST, CIFAR-10 and ImageNet dataset respectively. Moreover, to maintain a relatively small local Lipschitz constant of left half of the SPLITZ classifier, we set the threshold of clipped ReLU (see Sec B) as 1 for all datasets. For estimating the local Lipschitz constant of the left half of the classifier, the power iteration is 5, 5, 2 during the training for MNIST, CIFAR-10 and ImageNet respectively following from [19].